The expression that is equivalent to a3x+5 /a3-1

1
Expert's answer
2021-05-07T09:52:00-0400
a3x+5/a31=a3x+52=a3x+3a^{3x+5}/a^{3-1}=a^{3x+5-2}=a^{3x+3}